# Convert Language Model into a Value-based Strategic Planner

### v1

@@ 1 Introduction
Figure 1: Paradigm of straQ*.
In this paper, we propse a new framework called strategic Q* (straQ*), which converts LLM into a value-based strategic planner. We use the deep Q-learning (DQN) on LLM to provide a strategic Q function, with the strategy as the textual action. We use the averaged logits of actions to denote the Q value, and update the LLM parameter by the famous Bellman equation. By this manner, we convert the next-token prediction to next-strategy prediction, bootstrapping the TD loss of strategies instead of the original cross-entropy loss. This Q-net is used as a plug-and-play strategic planner, along with the conversation LLM produce the ultimate response. Our main contributions can be summarized as follows:
@@
(4) We design two reward machanisms including imitation and distillation, with the former better at automatic metrics, while the latter better at human scoring and generalization.
@@ 2 Preliminary
In this work, we formalize the ESC tasks as an Markov decision process (MDP). We define the MDP $\mathcal{M}$ at the strategy level with $\mathcal{M}=(\mathcal{S},\mathcal{A},\mathcal{T},\mathcal{R},\gamma)$. We define the action space $a_{t}\in\mathcal{A}$ is the set of possible strategies, and $\mathcal{T}:\mathcal{S}\times\mathcal{A}\rightarrow\mathcal{S}$ is the stepwise transition function.
@@
In value-based RL, the goal is to learn both the state-value function $V(s)$ and the state-action value function $Q(s,a)$, therefore determining the optimal strategy from $Q$ also achieves the highest expected discounted cumulative reward:
@@
which can be solved by the famous Bellman Equation:
$Q^{*}(s,a)=r(s,a)+\gamma\max_{a^{\prime}}Q^{*}(s^{\prime},a^{\prime}).$ (1)
in which the superscript indicates the next step.
Instead of explicitly implementing the above equation, Deep Q-learning (DQN) approximates the maximization of the right-hand side with the deep value networks:
@@
where $\mathcal{L}$ is the loss, $\theta$ and $\phi$ are parameters of the Q-net and the targe Q-net, respectively. $\phi$ can be periodically synchronized from $\theta$.
Figure 2: Framework of straQ*.
@@ 3.1 Task Definition
The problem of emotional-support conversation (ESC) can be characterized by an interleaved sequence of seeker $query$ and supporter $response$. To strengthen emotional-support performance, recent LM-based studies Liu et al. (2021) enhance the data content by augmenting the set of support strategies $\mathcal{A}$ and seeker emotions $\mathcal{E}$. For each conversation session, the background description can also be annotated on the session-level. Such augmented ESC can then be described as
@@
in which $T$ is the total number of conversation turn. At turn $t$, we denote the conversation history as
$h(t)=\{query,e(t),a(t),resp(t)\}_{0:T-1}$ (4)
Then the ESC sample at time $t$ can be alternatively expressed as $h(t),query(t),e(t),a(t),resp(t)$.
3.2 Framework
Our straQ* can be formally described as a 4-tuple $(\mathcal{S},\mathcal{A},\mathcal{R},\mathcal{T})$, where:
  • State. The state is a combination of description, emotion, history and query, i.e., $s=\{desc,e,h,query\}\in\mathcal{S}$.
  • Action. The strategy set is our action list.
  • Reward. An off-the-shelf foundation LLM generates the response upon on $a_{t}$, and the reward $r_{t}$ can be viewed as the instantaneous satisfaction of the seeker.
  • $\mathcal{T}:\mathcal{S}\times\mathcal{A}\rightarrow\mathcal{S}$ is the transition function. In the $t$-th turn, the seeker reacts further with a new query $q_{t+1}$ and emotion $e_{t+1}$. The history is updated by appending the current query and response, and the step is incremented by one.
@@ 3.3 Implementation on Language Models
Our implementation starts from a pretrained LLM, with the parameter of $\theta$. We assume there is an instruction template with the placehoder of $s$, denoted by $\mathcal{I}(s)$. This instruction can be concatenated with $a$, $\mathcal{I}\oplus a$. Both state and state-action values can be obtained from the semantic understanding of LLM:
$Q_{\theta}(s,a)$ $\leftarrow\text{LLM}_{\phi}(\mathcal{I}(s)\oplus a)$ (5)
where $\leftarrow$ means to average the action logits.
@@
By replacing the Q-net in Equation 2 by the above expressions, we finetune the LLM by the Bellman Equation loss on last token logit. We keep the setting of the target Q-net, which is the same LLM architecture, while its parameter $\phi$ are periodically synchronized from $\theta$.
As in standard language modeling, the causal masking of the transformer allows us to perform Bellman updates on entire sequences in parallel.
@@
Instead of decoding the next token, our finetuned LLM can produce the logits of available strategies, and the optimal strategy can be determined from the maximum logit
@@
3.4 Definition of Rewards
Definition of rewards may be crucial when the action sampling is constrained by an offline dataset. In this paper, we study two reward mechanisms:
(1) Distillation: for each ($s$, $a$) pair from the dataset, we let a strong-basis LLM (e.g., GPT-4) to provide a judge score from 0 to 5. The detailed score prompt is in the Appendix. Since this manner distills the knowledge from a teacher model, we call this variant as straQ*-distill.
(2) Imitation: we consider each ($s$, $a$) pair from the dataset is an expert demonstration, therefore, always assigning a $+1$ reward. To amplify the distribution, we randomly sample different $a$ and assigning a $-1$ reward. The postive-negative ratio is 1:1. Since this manner belongs to the imitation learning, this variant is called straQ*-imit.
Figure 3: Training loss curves of straQ*
@@ 3.4 Definition of Rewards
CoT 2.91 9.79 32.65 1.37
FSM 3.33 10.80 33.37 2.96
straQ*-distill (ours) 4.49 12.93 46.53 8.36
straQ*-imit (ours) 4.27 12.66 46.80 8.11
@@
4 Experiment
4.1 Setting
Implementation.
Llama3.2-1B-instruct AI@Meta (2024) is employed as the base model. Training is conducted on OpenRLHF Hu et al. (2024), with the learning rate of $65.0e-$ 6, window length of 2048 and epoch of 4. Batch size is set to 64. Experiments are conducted on 8 NVIDIA A100 80GB PCIe GPUs. In addition, the target network update frequency is set to 10, and the replay buffer size is 12,000.
Experiments.
Since our methodology is trained by annotated strategies, we employ ESconv as in-domain (ID) evaluation, and EmpatheticDialogues as the out-of-domain (OOD) evaluation, since EmpatheticDialogues does not have the strategy annotation. For the ID test, both strategy-related and response-related results can be provided. For the OOD test, we evaluate the zero-shot performance of the model finetuned on ESconv, with only response-related results provided.
Baselines.
We consider the following baselines:
(1) Direct: directly inference the LLM, with the same context.
(2) Direct-Refine: a straightforward refinement method in which the model revises its initial response to incorporate emotional support considerations.
(3) Self-Refine: a method (Madaan et al., 2023) initiates by generating feedback emphasizing emotional support from the initial response, then refining the response based on this feedback.
(4) CoT: uses the Chain-To-Thought prompt (Wei et al., 2022), which first generate the seeker’s emotion, which then guides the generation of strategy and response.
(5) FSM: the finite state machine Wang et al. (2024b) with finite sets of states and state-transitions triggered by inputs, and associated discrete actions.
4.2 Evaluation Methods
Automatic Metrics.
To evaluate the quality of strategy determination, we refer the evaluation methods proposed by Kang et al. (2024), which uses proficiency $\mathcal{Q}$ based on macro-F1 and the accuracy index (Acc) to represent the strategy accuracy, and preference bias $\mathcal{B}$ based on Bradley-Terry model (Bradley and Terry, 1952). Smaller $\mathcal{B}$ means less bias, therefore is better. For response quality, we utilize the famous Bleu-2 (B-2) and Rouge-L (R-L).
Human Scoring.
Similar with Kang et al. (2024), we annotate with the dimensions of Acceptance, Effectiveness, Sensitivity, Fluency, and Emotion, and the ultimate purpose, seeker’s Satisfaction. Scores are averaged over results of ESconv and EmpatheticDialogues.
@@ 4.3 Results
Figure 3 shows the training loss curve of straQ*for 500 steps (approximately 3 epochs). Although the loss fluctuates significantly with the steps, it converges obviously overall, and finally tends to be stable.
Automatic Metrics.
Table 1 presents the evaluation of automatic indicators in domain (ID) across different base reply models. Compared to CoT and FSM, straQ*achieves higher strategy accuracy, lower bias, and greater similarity to source dataset replies. While straQ*-imit shows higher Acc due to dataset-derived rewards, this does not imply superiority over straQ*-distill, as discussed later. Since the EmpatheticDialogues dataset lacks strategy labels, OOD results in Table 2 focus on dialogue content, where straQ*demonstrates advantages and strong generalization. Details of per-strategy results are shown in the appendix.
@@
The results of the crowdsourcing evaluation shown in Table 3 indicate that straQ*-distill outperforms the baseline methods in various metrics, such as Fluency, Emotion, and Satisfaction. It also performs better than the replies in the source data. Conversely, straQ*-imit is slightly lower than the source data in performance. This is because the reward signal of straQ*-distill is the GPT-4 score, which prevents the generation of obviously poor responses present in the dataset.
Figure 4: Distribution of strategy determined by straQ*.
Left: the confusion matrix. Right: the transition matrix.
4.4 Discussion
@@
Table 4 presents a statistical analysis of key performance indicators in reinforcement learning training. The <reward> represents GPT-4’s evaluation of generated replies on the test set, while the <value> reflects the model’s confidence in its decisions. Models trained with straQ*exhibit high <value>s, and straQ*-distill, leveraging GPT-4 scores as rewards, achieves a higher <reward>.
@@
straQ*-distill 3.99 424.78
straQ*-imit 3.72 445.95
@@
Figure 5: Sensitivity plots of straQ* on different $\gamma$ and model sizes. Metrics include Acc, $\mathcal{Q}$, B2 and R-L.
@@
The confusion matrix shown in Figure 4 indicates that for most strategies, the color is the darkest along the main diagonal. In the transition matrix, the color in the upper right triangular area is significantly darker than that in the lower left area, which is consistent with the three-stage theory in ESC.
@@
supporter - Hi! Hope you are doing well? 4 -
@@
supporter - How can I help you today? 3 -
@@
supporter Res.& Par. So as far as I understand that you had issues with your wife due to your drinking. Am I right? 3 7.6
straQ*-distill Res.& Par. It sounds like your drinking habits have been causing difficulties in your marriage for a while. 3 10.1
Direct Res.& Par. So, it sounds like there’s been a bit of tension in your marriage lately due to your drinking habits, and it even led to a confrontation that’s caused a rift between you and your wife. Can you tell me more about what’s been going on with your drinking? 4 8.0
@@
supporter Aff.& Rea. That is really a serious problem. I know it must be very tough time for you. Hope you are trying your best to save your marriage. 2 5.4
straQ*-distill Ref. That must be really difficult for you. I can imagine how much stress and worry this situation has brought into your life. 4 8.3
Direct Res.& Par. Just to make sure I understand correctly, it sounds like there was a confrontation with your wife about your drinking habits, and now you two aren’t speaking. Is that a fair summary of what happened? 3 4.7
@@
supporter Ref. That’s is a great initiative. I know it is tough for you but not impossible. I can feel how hard you are trying t save your relationship. I appreciate you for that. 4 4.0
straQ*-distill Aff.& Rea. That’s a great step forward! It takes a lot of strength and commitment to make such a change, and you should be proud of yourself. 5 5.0
Direct Res.& Par. Just to make sure I understand correctly, you’ve been trying to cut down on your drinking to address your wife’s concerns, but despite your efforts, the issue still seems to be causing tension in your marriage, and now you’re not talking to each other. Is that a correct summary? 2 2.0
@@
4.5 Ablation Study
We conduct the following ablation studies:
  • w/ value head: Model is connected to a fully-connected network, and the output layer has only one neuron serving as the strategy value.
  • auto-regressive: Cross-entropy loss is used to calculate the loss instead of the reward signal.
Table 5 shows the straQ*model demonstrates better performance in various automatic metrics, which reflects the effectiveness of the proposed method.
4.6 Sensitivity Analysis
Figure 5 shows when $\gamma$ is within a certain range (0.8 to 0.9), its impact on the final performance is relatively small. However, when $\gamma$ is 0.99, the performance of the model decreases significantly. In addition, as the model size increases from 1B to 8B, the model performance also rises slowly. Nevertheless, considering the comprehensive factors of inference speed and computing resources, a 1B model is already enough.
4.7 Case
Table 6 presents the response scenarios of different models to the same set of multi-turn dialogues. straQ*-distill effectively achieves the three-stage strategy transfer and generally attains higher scores. In contrast, the strategy of the direct group remains unchanged throughout the multi-turn dialogues. As the number of turns increases, it becomes increasingly inappropriate, leading to lower scores.
5 Related Work
Typically, strategy learning occurs in goal-oriented conversations. Representative examples include DAT Li et al. (2024b). It represents conversations in a structured way by defining dialogue action tags and a multi-turn planner, and then generates responses based on these tags and pre-set planning strategies. In addition, DQ-HGAN Li et al. (2024a) realizes emotional support through the Graph Attention Network. ArCHer Zhou et al. (2024) proposes a hierarchical multi-turn RL algorithm to improve the efficiency and performance of language models in multi-turn goal-oriented decision-making tasks, and Q-star Wang et al. (2024a) improves multi-step reasoning for LLMs. However, such works usually do not have a clear explicit strategy, and their evaluation often relies on ground truth values.
In contrast, our straQ* learns to plan multi-turn explicit strategies offline from scoring feedback, and is decoupled from the dialogue model, which facilitates training and deployment..
6 Conclusion
In this paper, based on Q-learning, we propose a method named straQ*that can provide long-context emotional support conversation in multi-turn dialogue scenarios. We train to obtain the straQ*-imit model, which has a higher strategy accuracy and a higher response consistency on the Esconv dataset. In addition, we proposes the straQ*-distill method, which is trained from the reward signals of GPT-4’s scoring and achieves more excellent human evaluation results, and can be generalized to a broader range of scenarios.
7 Limitation
Appendix A Further Implementation Details
A.1 Dataset details
Emotional Support Conversations.
Emotional Support Conversation (ESC) is a task aimed at alleviating users’ negative emotions (e.g., anxiety, depression), where supporters assist seekers in managing emotions triggered by issues like work crises or interpersonal conflicts. Unlike emotion recognition tasks, ESC integrates psychological counseling mechanisms into the dialogue generation process, offering a deeper, context-sensitive solution for emotion regulation. In the practice, ESconv have exactly the aforementioned architecture while EmpatheticDialogues lacks the Strategy.
ESconv.
Motivated by the Helping Skills Theory Hill (2009), Liu et al. (2021) divides ESC into three sequential stages: Exploration, Comforting, and Action, and proposes a dataset called ESconv. For each sample, the conversation is multi-turn, with the dialogue background and user emotion annotated. Upon each utterance of the supporter, 8 distinct support strategies are annotated. Table LABEL:tab:ex_ESConv exhibits an example of ESConv.
      Strategies Abbr. Definitions
Table 7: Strategy definitions and Abbreviations of ESConv.
Emotion Type # Occurrence
anger 111
anxiety 354
depression 334
disgust 40
fear 95
nervousness 13
sadness 308
shame 42
Table 8: Emotion statistics of ESConv.
Table 7 provides definitions of support strategies in ESConv. Table 8 lists the emotion types and their occurrences in the dataset. The emotion types include anger, anxiety, depression, disgust, fear, nervousness, sadness, and shame. Table 10 summarizes the basic statistical information of ESConv.
EmpatheticDialogues.
EmpatheticDialogues is a dataset that consists of empathetic conversations. It aims to help in the development of empathetic language models by providing a large number of dialogues that express empathy.
A.2 Prompt format
To further strenthen the understanding capability of LLM on the strategy selection, we define the instruction as a multi-choice question (MCQ), forcing the LLM to choose one of the option numbers, instead of a plain question. Below is the content of the instruction template $\mathcal{I}$:
Emotion: {$e$}
Description: {$desp$}
{$h$}
The seeker’s current query is:
{$query$}
strategy #(1) {$a_{1}$}
strategy #(k) {$a_{k}$}
The prompt used in response generation:
Emotion: {$e$}
Description: {$desp$}
{$h$}
The seeker’s current query is:
{$query$}
The current response strategy is:
{$a$}
template of GPT-4 scoring.
Emotion: {$e$}
Description: {$desp$}
{$h$}
The seeker’s current query is:
{$query$}
Please evaluate whether the response is appropriate:
{$resp$}
@@ A.2 Prompt format
The performance of the model often differs among various strategies. For humans, straightforward strategies like "Question" and "Providing Suggestions" may be more convenient to learn and apply. As a result, these strategies also appear relatively more frequently in the source data. Table 9 shows that for both the SFT and straQ*models, the above-mentioned strategies also have higher values of Acc, $\mathcal{Q}$, and R-L value. Moreover, compared with the SFT model, in addition to having higher indicators such as the average Acc, the straQ*model is more balanced among different strategies and avoids the occurrence of zero values.

### v2

@@ 1 Introduction
Figure 1: Paradigm of straQ*. A plug-and-play LLM-based planner selects the optimal strategy from maximized $Q$, then steers the LLM to enhance the response.
In this paper, we propose a new framework called strategic Q* (straQ*), which converts LLM into a value-based strategic planner. We use the deep Q-learning (DQN) on LLM to provide a strategic Q function, with the strategy as the textual action. We use the averaged logits of actions to denote the Q value, and update the LLM parameter by the famous Bellman equation. By this manner, we convert the next-token prediction to next-strategy prediction, bootstrapping the TD loss of strategies instead of the original cross-entropy loss. This Q-net is used as a plug-and-play strategic planner, along with the conversation LLM to produce the ultimate response. Our main contributions can be summarized as follows:
@@
(4) We design two reward mechanisms including imitation and distillation, with the former better at automatic metrics, while the latter better at human scoring and generalization.
@@ 2 Preliminary
The Markov decision process (MDP) is usually defined as a 5-tuple $(\mathcal{S},\mathcal{A},\mathcal{R},\mathcal{T},\gamma)$, where $\mathcal{S}$ is the state set, $\mathcal{A}$ is the action set, $\mathcal{R}$ is the reward set, $\gamma$ is the discounting factor of rewards, and $\mathcal{T}:\mathcal{S}\times\mathcal{A}\rightarrow\mathcal{S}$ is the state transition function. In this work, we formalize the ESC task as a strategy-level MDP, with the action space $\mathcal{A}=\{a\}$ as the set of possible strategies.
@@
In value-based RL, the goal is to learn the state-value function $V(s)$ or the state-action value function $Q(s,a)$, such that the determined action achieves the highest expected discounted cumulative reward:
@@
which is solved by the famous Bellman Equation:
$Q^{*}(s,a)=r(s,a)+\gamma\max_{a^{\prime}}Q^{*}(s^{\prime},a^{\prime})$ (1)
in which the superscript indicates the next step. Instead of explicitly implementing the above equation, Deep Q-learning (DQN) approximates the maximization of the right-hand side with the deep value networks:
@@
where $\mathcal{L}$ is the loss, $\theta$ and $\phi$ are parameters of the Q-net and the target Q-net, respectively. $\phi$ can be periodically synchronized from $\theta$.
Figure 2: The training framework of straQ*. Averaged log probability of action tokens is defined as $Q(s,a)$ which deduces the training loss from Bellman Equation.
@@ 3.1 Task Definition
The problem of emotional-support conversation (ESC) can be characterized by an interleaved sequence of seeker $query$ and supporter $response$. To strengthen emotional-support performance, recent studies Liu et al. (2021); Rashkin et al. (2019) enhance the data content by augmenting the set of support strategies $\mathcal{A}$ and seeker emotions $\mathcal{E}$. For each conversation session, the background $description$ is also annotated on the session-level. Such augmented ESC can then be described as
@@
in which $desp$ and $resp$ are the abbreviations of $descrition$ and $response$, and $T$ is the total number of conversation turns. At turn $t$, we denote the conversation history as
$h(t)=\{query(t),e(t),a(t),resp(t)\}_{0:T-1}$ (4)
Then the ESC sample at time $t$ can be alternatively expressed as $\{h(t),query(t),e(t),a(t),resp(t)\}$.
3.2 System Variables
We define important system variables as follows:
$\bullet$ State: The state is a combination of description, emotion, history and query, i.e., $s=\{desc,e,h,query\}\in\mathcal{S}$.
$\bullet$ Action: The conversational strategy, $a\in\mathcal{A}$.
$\bullet$ Reward: The reward $r_{t}$ can be viewed as the instantaneous satisfaction of the seeker, which can be either inferred from an annotated datasets, or generated by an off-the-shelf model evaluator (LLM-as-the-Judge).
$\bullet$ $\mathcal{T}:\mathcal{S}\times\mathcal{A}\rightarrow\mathcal{S}$ is the transition function. After the $t$-th turn, $h$ is updated by appending the current $query(t)$ and $resp(t)$, the seeker reacts further with new $query(t+1)$ and $e(t+1)$, and the step is incremented by one.
@@ 3.3 Implementation on Language Models
Our implementation starts from a pretrained LLM, with the parameter of $\theta$. We assume there is an instruction template with the placeholder of $s$, denoted by $\mathcal{I}(s)$. This instruction can be concatenated with $a$, $\mathcal{I}\oplus a$. Both state and state-action values can be obtained from the semantic understanding of LLM:
$Q_{\theta}(s,a)$ $\leftarrowtail\text{LLM}_{\theta}(\mathcal{I}(s)\oplus a)$ (5)
[…]
@@
By replacing the Q-net in Equation 2 by the above expressions, we finetune the LLM by the Bellman Equation loss on last token logit. As in standard language modeling, the causal masking of the transformer allows us to perform Bellman updates on entire sequences in parallel. Figure 2 exhibits this training framework.
We keep the setting of the target Q-net, which is the same LLM architecture, while its parameter $\phi$ are periodically synchronized from $\theta$.
@@
Instead of decoding the next token, the finetuned LLM produces logits of available strategies, and the optimal strategy can be determined from the maximum logit
@@
Figure 3: Diagram of our pipeline. Reward is annotated by the imitation or distillation strategy. The action is chosen from the strategy which maximize the $Q$ value.
Instruction template.
We briefly exhibit our instruction $\mathcal{I}(s)$ here:
[…]
with the full version in Appendix B. By forming the prompt To further strengthen the understanding capability of LLM on the strategy selection, we formulate $\mathcal{I}$ as a multi-choice question (MCQ), instead of a plain question, forcing the LLM to choose one of the option numbers. Accordingly, the action set becomes the set of possible strategy index $a\in\mathcal{A}:=\{1,2,\cdots,K\}$ where $K$ is the total number of strategies.
3.4 Reward Definitions
Choice of rewards may be crucial especially when the sampling is constrained by an offline dataset. In this paper, we study two reward mechanisms:
(1) Distillation: for each ($s$, $a$) pair from the dataset, we let a strong-basis LLM (e.g., GPT-4) to provide a judge score from 0 to 5 (The detailed judge prompt is in Appendix B). Since this manner distills the knowledge from a teacher model, we call this variant as straQ*-distill.
(2) Imitation: we consider each ($s$, $a$) pair from the dataset is an expert demonstration, therefore, always assigned with $r$ of $+1$. To amplify the distribution, we randomly sample a different $a$ and assign with $r$ of $-1$. The positive-negative ratio is 1:1. Since this manner imitates the positive samples directly, this variant is called straQ*-imit.
Figure 3 shows the entire pipeline of straQ*.
Figure 4: Training loss curve of straQ*.
Strategies Abbr. Stage
Question Que. I
Restatement or Paraphrasing Res.& Par. I
Reflection of Feelings Ref. II
Self-disclosure Self-Dis. II
Affirmation and Reassurance Aff.& Rea. III
Providing Suggestions Pro. III
Information Inf. III
Others Others -
Table 1: Strategy names, abbreviations and stages.
4 Experiment
4.1 Setting
Implementation.
Llama3.2-1B-instruct AI@Meta (2024) is employed as the base model. Training is conducted on OpenRLHF Hu et al. (2024), with the learning rate of $65.0e-$ 6, window length of 2048, batch size of 64, and epoch of 4. The target network update frequency is set to 10, the replay buffer size is 12,000, and $\gamma=0.85$.
Datasets.
Training of straQ* requires the annotation of strategies. We use ESConv Liu et al. (2021) as the training set and also the in-domain (ID) test. ESConv provides $K=8$ strategies, which belong to three ESC stages: Exploration (I), Comforting (II) and Action (III). Table 1 shows their full names, abbreviations and corresponding stages.
Furthermore, EmpatheticDialogues Rashkin et al. (2019) is employed as the out-of-domain (OOD) evaluation, since EmpatheticDialogues does not have the strategy annotation. For the ID test, both strategy-related and response-related results can be provided. For the OOD test, only zero-shot response-related results are provided. Appendix A provides a more detailed introduction of ESConv and EmpatheticDialogues.
4.2 Evaluation Methods
Automatic Metrics.
To evaluate the quality of strategy determination, we refer the evaluation methods proposed by Kang et al. (2024), which uses proficiency $\mathcal{Q}$ based on macro-F1, and preference bias $\mathcal{B}$ based on Bradley-Terry model (Bradley and Terry, 1952). Smaller $\mathcal{B}$ means less bias, therefore is better. We also include the strategy prediction accuracy (Acc). For response quality, we utilize the famous Bleu-2 (B-2), Rouge-L (R-L), Distinct-2 (D-2) and CIDEr, calculating from the similarity with the ground truth response.
Human Scoring.
Similar with Kang et al. (2024), we annotate with the dimensions of Acceptance, Effectiveness, Sensitivity, Fluency, and Emotion, and the ultimate purpose, seeker’s Satisfaction.
Baselines.
We consider the following baselines:
(1) Direct: directly inference the LLM.
(2) Direct-Refine: the model immediately revises its response within the same utterance to incorporate emotional support considerations.
(3) Self-Refine (Madaan et al., 2023): the model considers the emotional support, generates a feedback from the initial response, then refines the response based on the feedback.
(4) CoT (Wei et al., 2022): steered by the chain-of-thought prompt, the model first identifies emotion, then generates strategy, and finally response.
(5) FSM Wang et al. (2024b): the finite state machine with finite sets of states and state-transitions triggered by inputs, and associated discrete actions.
@@ 4.2 Evaluation Methods
+ CoT 2.91 9.79 32.65 1.37
+ FSM 3.33 10.80 33.37 2.96
+ 1B straQ*-distill (ours) 4.49 12.93 46.53 8.36
+ 1B straQ*-imit (ours) 4.27 12.66 46.80 8.11
@@ 4.3 Results
Figure 4 shows the training loss curve of straQ*for 500 steps (approximately 3 epochs). Although the loss initially fluctuates significantly, it adapts to the new training paradigm, and finally tends to be stable.
Automatic Evaluations.
Table 2 presents the automatic metrics on the ID evaluation, with the basis of either the original LLM, or the specifically finetuned version. Compared to baselines, straQ*generally achieves higher strategy accuracy, lower bias, and higher similarity to the ground truth responses. Furthermore, straQ*-imit performs better than straQ*-distill on this setting, suggesting that the imitation-version of rewards result in better ID performance.
In Table 3, we further compare the OOD results of the models finetuned by ESConv, with the strategy lists inferred from ESConv. Results suggest that straQ*demonstrates strong generalization than these baselines. Specifically, straQ*-distill surpasses straQ*-imit this time, indicating the distilled knowledge from the teacher model is more general than simply imitating a limited dataset.
@@
The results of the crowdsourcing evaluation shown in Table 4 indicate that straQ*-distill outperforms the baseline methods in various metrics, such as Fluency, Emotion, and Satisfaction. It also performs better than the replies in the source data. Conversely, straQ*-imit is slightly lower than the source data in performance. Using the GPT-4 score as reward, straQ* can determine strategies more from the aspect of performance optimization, not simply imitating the demonstration.
Figure 5: Distribution of strategy determined by straQ*. Strategies are labeled with the stage index (I, II, III) which represents the general scenario Exploration $\rightarrow$ Comforting $\rightarrow$ Action in ESC. Left: the confusion matrix (acted strategy (row) VS ground truth strategy (column)). Right: the transition matrix (acted strategy (row) VS the next-acted strategy (column)).
Ablation Study.
Two ablations are studied:
(1) w/ value head: append the model with a classification head which produces the score logit.
(2) auto-regressive: keep the cross-entropy loss with the ground truth action as the target text.
In more detail, w/ value head is the usual solution for a reward model in RLHF, while auto-regressive can be viewed as a standard fine-tuning solution for the strategic planner. Table 6 shows that straQ*outperforms both of them in various automatic metrics, indicating our methodology can better align with the strategy semantics and more accurately capture the strategic value.
Sensitivity Analysis.
Figure 6 shows the ID performance evolutions on different $\gamma$ choices. Smaller $\gamma$ means we are more focused on the transient performance and relatively neglect the long-term value. Results show that the optimal accuracy of strategy happens on $\gamma=0.9$, while the best response-related metrics correspond to $\gamma=0.85$. Because B-2 and R-L are similarity-based, the current reward is more relative to them than future rewards. Therefore, this observation is reasonable.
4.4 Discussions
Scalability and application.
Figure 6 (bottom-right) also compares the B-2 results on different model sizes. As the model becomes larger, the performance also increases, indicating straQ* can have good scalability. However, larger models result in higher computation overhead and slower speed, which may hinder the practical application of straQ*. Therefore, in the formal application, we still adhere to the 1B choice, employing it as a lightweight planner.
From previous results, we utilize straQ*-distill in the actual application to have better generalization and better alignment with human knowledge.
@@ 4.4 Discussions
Table 5 further analyzes two important indicators of value-based RL, the averaged rewards and values. In this analysis, the rewards are provided by GPT-4. straQ* achieve both higher <reward> and <value> than direct inference of the base model, as well as the annotation of original dataset. This result shows that straQ* statistically obtains higher returns, which is the primary purpose of Q-Learning.
@@
straQ*-distill (ours) 3.99 424.78
straQ*-imit (ours) 3.72 445.95
@@
Figure 6: Sensitivity plots of straQ*-imit on different $\gamma$ and model sizes. Metrics include Acc, $\mathcal{Q}$, B2, and R-L.
@@
Figure 5 (Left) exhibits the confusion matrix of strategies, with the rows representing the prediction, and the columns representing the ground truth. Results show that most occurrences happen on the diagonal grids, verifying the prediction accuracy.
Figure 5 (Right) visualizes the transition matrix. A grid $(i,j)$ means the strategy $i$ to the strategy $j$, where the strategies are sorted from their ESC stages (from I to III) for both rows and columns. Therefore, transitions from an early stage to a later stage should occur on the upper-triangle region of the transition matrix. Results in the figure validate this proposition.
@@
supporter - Hi! Hope you are doing well? - -
@@
supporter - How can I help you today? - -
@@
supporter Res.& Par. (I) So as far as I understand that you had issues with your wife due to your drinking. Am I right? 3 7.6
Direct Res.& Par. (I) So, it sounds like there’s been a bit of tension in your marriage lately due to your drinking habits, and it even led to a confrontation that’s caused a rift between you and your wife. Can you tell me more about what’s been going on with your drinking? 4 8.0
straQ*-distill Res.& Par. (I) It sounds like your drinking habits have been causing difficulties in your marriage for a while. 3 10.1
@@
supporter Aff.& Rea. (II) That is really a serious problem. I know it must be very tough time for you. Hope you are trying your best to save your marriage. 2 5.4
Direct Res.& Par. (I) Just to make sure I understand correctly, it sounds like there was a confrontation with your wife about your drinking habits, and now you two aren’t speaking. Is that a fair summary of what happened? 3 4.7
straQ*-distill Ref. (II) That must be really difficult for you. I can imagine how much stress and worry this situation has brought into your life. 4 8.3
@@
supporter Ref. (II) That’s is a great initiative. I know it is tough for you but not impossible. I can feel how hard you are trying t save your relationship. I appreciate you for that. 4 4.0
Direct Res.& Par. (I) Just to make sure I understand correctly, you’ve been trying to cut down on your drinking to address your wife’s concerns, but despite your efforts, the issue still seems to be causing tension in your marriage, and now you’re not talking to each other. Is that a correct summary? 2 2.0
straQ*-distill Aff.& Rea. (III) That’s a great step forward! It takes a lot of strength and commitment to make such a change, and you should be proud of yourself. 5 5.0
@@
Strategies’ popularities and occurrences may differ in nature. For example, straightforward strategies like "Question" and "Providing Suggestions" may be more convenient to learn and apply. A reasonable strategic planner should reflect this frequency difference, but also avoid significant bias (e.g., always determines the most frequent strategy). To further study the strategy-related performance, we further exhibit the per-strategy results in Table 8, with the comparison between straQ*-distill and SFT. One can observe that straQ* model is more balanced across different strategies, and SFT even have zero predictions on the strategy of "Information".